Suppose E = 6.0 V. Find the current through resistor c and find the potential difference across resistor c.

5.0 ? 5.0 ? ?
Answer #1

here,

(Rc + Rd ) are in parallel with Rb, their equivalent resistance , R1 = Rb * ( Rc + Rd) /( Rb + Rc + Rd)

R1 = ( 5 + 5) * ( 10 ) /(10 + 5 + 5)

R1 = 5 ohm

R1 and Ra are in series , their equivalent resistance , Req = R1 + Ra = 10 ohm

the current through a , I = E/Req

I = 6 /10 A = 0.6 A

the potential difference across a , Va = I * Ra

Va = 3 V


answered by: ANURANJAN SARSAM